OpenStudy (anonymous):

medal and fan first help plz What is the quotient of three and one over three divided by five over seven in simplest form? three and five over twenty-one three and seven over fifteen four and two over three four and eight over twenty-one

OpenStudy (anonymous):

Is that the expression you're supposed to simplify? \[\frac{ 3+1 }{ \frac{ 3 }{ \frac{ 5 }{ 7 } } }\]

OpenStudy (anonymous):

If so, you can aleady add what's in the numenator together: 3+1 = 4 Dividing by a fraction is the same as multiplying with the turned over fraction, like this \[4.\frac{ \frac{ 5 }{ 7 } }{ 3 }\] We can now rearrange the fractions as \[\frac{ 4 }{ 3 }.\frac{ 5 }{ 7 }\] Now we calculate everything and get \[\frac{ 20 }{ 21 }\]

OpenStudy (anonymous):

thats not one of the answers but thankyou

OpenStudy (anonymous):

I know, but maybe I misunderstood what was meant with over and so on.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

f

OpenStudy (anonymous):

But maybe I have given you an idea what to play around with.